susan has 10 pockets and 44 dollars bills. she wants to arrange the money so that there are a different number of dollars in each pocket. can she do it? explain.
Found 2 solutions by Alan3354, Theo:
Answer by Alan3354(69443)

Assuming she can't use IOU's or tear bills in pieces:
0 is the minimum # of bills
---
The sum of 0 to 9 is 45.
Can't be done.

Answer by Theo(13342)

i don't think so.
the smallest amount she can put in each pocket adds up to 45.
this assumes that one of the pockets is left empty.
0 + 1 + 2 + 3 + 4 + 5 + 6 + 7 + 8 + 9 + 10 = 45.
she couldn't leave 2 pockets empty because then there would not be a different number of dollars in each pocket.
with only 44 dollars, two of the pockets would have to have the same amount.
just subtract 1 from any of the pockets except the one that has 0, and you get a total of 44 with two of the pockets having the same amount.
for example: 10 - 9 = 9 but there is already a pocket with 9.
5 - 1 = 4 but there is already a pocket with 4.
etc.....